OpenStudy (anonymous):

The vertex if the parabola below is at the point (3,-3) and the point (0,-2) is on the parabola . Write the equation of the parabola.

OpenStudy (anonymous):

If we know the vertex, then, that means: \[ y=a(x-x_0)^2+y_0\implies\\ y=a(x-3)^2-3 \]And, we plug in values to find \(a\): \[ \begin{align} -2&=a(0-3)^2-3\implies\\ -2&=9a-3\implies\\ 1&=9a\implies\\ a&=\frac{1}{9} \end{align} \]Therefore, our formula is: \[ y=\frac{1}{9}(x-3)^2-3 \]

OpenStudy (anonymous):

Thank you <3